# EMMA: Scaling Mobile Manipulation via Egocentric Human Data

### v2

@@ Abstract
Scaling mobile manipulation imitation learning is bottlenecked by expensive mobile robot teleoperation. We present Egocentric Mobile MAnipulation (EMMA), an end-to-end framework training mobile manipulation policies from human mobile manipulation data with static robot data, sidestepping mobile teleoperation. To accomplish this, we co-train human full-body motion data with static robot data. In our experiments across three real-world tasks, EMMA demonstrates comparable performance to baselines trained on teleoperated mobile robot data (Mobile ALOHA), achieving higher or equivalent task performance in full task success. We find that EMMA is able to generalize to new spatial configurations and scenes, and we observe positive performance scaling as we increase the hours of human data, opening new avenues for scalable robotic learning in real-world environments. Details of this project can be found at: https://ego-moma.github.io/
@@ I Introduction
Concurrently, recent work in cross-embodiment learning has explored human data as a scalable source of data for training end-to-end robot policies. Critically, human video data is cheap to collect, does not require a physical robot, and can be collected ergonomically via XR wearables [egomimic, qiu2025humanoidpolicyhuman]. Human video data has been leveraged to train better visual representations [nair2022r3muniversalvisualrepresentation], extract high-level object affordance [wang2023mimicplay, xu2024flowcrossdomainmanipulationinterface, wen2024anypointtrajectorymodelingpolicy, bharadhwaj2024track2actpredictingpointtracks], or even as direct action supervision via co-training [egomimic, qiu2025humanoidpolicyhuman]. However, these works have predominantly focused on table-top manipulation.
To this end, we introduce Egocentric Mobile Manipulation (EMMA), an end-to-end system to train mobile manipulation policies solely from robot static-manipulation and human mobile-manipulation data. While the robot data is captured via teleoperation, the human data is captured simply by wearing Project Aria glasses [engel2023projectarianewtool], thus avoiding costly mobile manipulation teleoperation. EMMA represents a step toward a new data paradigm in robot learning, where we imbue a robot with new skills by combining a set of robot teleoperation data with a more diverse pool of human demonstrations. In this work, our robot teleoperation data contains no mobile manipulation demonstrations, and we show that this skill can be effectively transferred from human data.
EMMA presents a full-stack framework consisting of 1) an action retargeting pipeline which translates human mobile manipulation actions to a bimanual robot with a differential-drive base, 2) a unified architecture designed for co-training on heterogeneous human and robot data, and 3) an auxiliary phase identification mechanism that modulates between navigation and manipulation modes during inference. Therefore, it can switch control modes between navigation and manipulation phases based on predictions, preventing unintended base drift during manipulation and out-of-distribution arm movements during navigation.
We evaluate EMMA across four real-world mobile manipulation tasks: table service, handover wine, grocery shopping and push chair. Our experiments reveal three key findings. First, EMMA achieves superior performance compared to baselines trained on teleoperated mobile robot data, showing that human demonstrations can replace costly mobile teleoperation. Second, we observe favorable scaling properties—each additional portion of human data yields greater performance than an equivalent portion of teleoperated robot data. Third, EMMA exhibits robust generalization performance, successfully transferring navigation and manipulation skills to novel environments seen only in human demonstrations. These results open new avenues for scalable robotic learning in real-world environments.
@@ II Related Work
Learning for Mobile Manipulation. Building on successes in static manipulation, recent works have explored learning-based approaches for mobile manipulation that include skill primitives [sun2022fully, wu2023m, wu2023tidybot, kuang2025skillblender], reinforcement learning with decomposed action spaces  [gu2022multi, yokoyama2023asc, xia2021relmogen, xiong2024adaptive], and whole-body control objectives [fu2023deep, yang2024harmonic, hu2023causal, liu2024opt2skill, ben2025homie, wang2024hypermotion]. Unlike these approaches, end-to-end imitation learning enables mapping raw pixel information to whole-body actions, showing promising results through large-scale training [ahn2022can, brohan2022rt, shafiullah2023bringing, pi0.5, kim24openvla, moo2023arxiv]. A critical challenge is to collect high-quality mobile manipulation demonstrations. Pioneering works [jiang2025behavior, fu2024humanplus] have developed new teleoperation systems to facilitate data collection, including tethered leader-follower system [fu2024humanplus], VR headsets [chen2024arcapcollectinghighqualityhuman, lu2025mobiletelevisionpredictivemotionpriors], motion-capture suits [cisneros2022team, dafarra2024icub3, wholebodygeomretarget], smartphone-based control [pmlr-v164-wong22a], kinesthetic teaching [MOMO-FORCE], and full-body teleoperation for humanoid robots [fu2024humanplus]. However, despite these advances, collecting high-quality mobile manipulation data for diverse scenarios at scale remains a challenge. We propose to leverage egocentric human data (data collected with first-person perspective) captured by wearable devices as an alternative for scaling up imitation learning for mobile manipulation.
Robot Learning from Human Data. Recent work has focused on two complementary themes—leveraging human videos to bootstrap robot learning and finetuning with reinforcement learning for robust policies. In manipulation, co-training on paired egocentric human and robot demonstrations has been shown to boost skill performance [egomimic], while zero- or few-shot transfer from human videos can be enabled by image inpainting and motion-track priors [lepert2025phantomtrainingrobotsrobots, ren2025motion]. Building on these, hierarchical planners extract latent action sequences from humans and distill them via retargeters into whole-body controllers [wang2023mimicplay, fu2024humanplus, ImitationNet]. In navigation, fusing imitation with RL—by turning keypoint matches into rewards or bootstrapping from behavior cloning—yields more reliable policies  [karnan2022voila, ramrakhya2023pirlnav], and inverse-dynamics models can pseudo-label passive egocentric video to distill intent-conditioned affordance subroutines [navsubroutine_egocentricvideos]. We aim to train mobile manipulation policies from human data in a unified learning framework.
@@ III Hardware and Data Preliminaries
Human and Robot Data Streams. We collect distinct data streams from both human demonstrations ($\mathcal{D}_{H}$) and robot execution ($\mathcal{D}_{R}$). A key shared data stream is the egocentric RGB image $I_{\text{ego}}$, generated by the Aria glasses worn by the human or mounted on the robot. The robot provides RGB streams from its two wrist cameras, $I_{\text{wrist}}$. Proprioceptive data streams capture the state of each embodiment: For the human, we leverage the Aria Machine Perception Service (MPS) [aria_mps] to estimate bimanual 3D hand poses $H_{p}\in SE(3)\times SE(3)$ and the 3D head pose $H_{d}\in SE(3)$ (in a SLAM-based world frame). For the robot, proprioception includes the state of its bimanual arms via joint positions $R_{q}\in\mathbb{R}^{2\times 7}$ (including gripper state) and the corresponding end-effector poses $R_{p}\in SE(3)\times SE(3)$. These raw streams serve as the inputs to our system; their processing and transformation into policy inputs and action labels is a core algorithmic contribution detailed in Sec. IV-A.
@@ IV EMMA: System and Algorithm
EMMA is a scalable full-stack system, which enables (1) direct transfer of navigation skills from egocentric human data to a differential-drive mobile manipulator and (2) scaling up full mobile manipulation policy performance by co-training on both human mobile manipulation data and robot static manipulation data.
@@ IV-B Human and Robot Data Co-training
Following the data processing and retargeting detailed in Sec. IV-A, which formally bridges the kinematic and distributional gaps, the goal of our system is to transfer knowledge from egocentric human mobile manipulation data ($D_{H}$) to scale policy performance, while still leveraging limited static robot manipulation data ($D_{R}$) for precise manipulation steps. But large domain gaps still exist in the two data sources, in both sensing modalities and distributions. Thus, effectively learning from these data necessitates a unified learning framework capable of processing heterogeneous data sources.
Inspired by recent works in cross-embodiment policy learning [liu2024rdt, wang2024hpt, ghosh2024octo], we design an architecture based on a decoder-only Transformer with modality-specific input stems, a shared trunk, and multiple action and auxiliary output heads (Fig. 2). Our goal is to transfer semantics of navigation and manipulation phases from human data. In addition, we want both the manipulation action prediction and navigation action prediction to benefit from large-scale human data. This motivates an architecture that shares a majority of the learnable parameters between the human and robot embodiments to promote shared representation, inspired by EgoMimic [egomimic].
@@
Co-training from heterogeneous sources. The model is trained jointly on batches drawn from two sources: (1) the collected human mobile manipulation demonstrations $D_{H}$ (with navigation actions processed by the retargeting module described in Sec. IV-A) and (2) static robot manipulation demonstrations $D_{R}$ (e.g., collected via teleoperation). When processing a human data batch, the human proprioception, shared ego vision stem, human manipulation action head, navigation action head and the shared phase prediction head are active. For a robot batch, the robot proprioception, wrist image stem, shared ego vision stem, and robot manipulation action head are active. The shared trunk and ego vision stem is updated by all sources and modalities, forcing them learn versatile representations. The navigation head primarily learns from the retargeted human data, transferring human navigation strategies to the robot. The complete architecture is illustrated in Fig. 2
@@ IV-C Auxiliary Phase Identification and Control Modulation
Phase Detection. We identify phases based on motion dynamics. For each frame, we compute the ratio of hand velocity to head velocity. Manipulation phases exhibit high hand-to-head velocity ratios ($\tau_{\text{ratio}}=2.0$) with low absolute head velocity ($\tau_{\text{head}}=0.4m/s$), indicating dexterous hand motion while stationary. We filter candidate segments shorter than $\tau_{\text{duration}}=30$ frames to remove transient gestures.
@@
To spatially localize $K$ manipulation zones, we fit a Gaussian Mixture Model (GMM) with $K=2$ components to head positions during these high-ratio periods. Each frame is classified as manipulation (phase 0) if its probability density under the GMM exceeds $\tau_{pdf}$, otherwise navigation (phase 1). GMM provides direct probability density estimates needed for phase classification without requiring labeled training data, while being computationally efficient and interpretable for spatial clustering. To assess accuracy, we hand-annotate 10% of uniformly sampled continuous segments from four tasks that require phase identification (see Section V). We report mean over frames (MoF), defined as the number of frames where the manual label matches the dataset label divided by the total number of sampled frames.
@@
This phase-aware modulation removes the base action noise introduced from head movements during manipulation and ensures smooth switch between the two phases.
@@ V Experiments
We aim to validate three key hypotheses. H1: EMMA can achieve performance comparable to systems trained on teleoperated mobile manipulation data. H2: Key design decisions of EMMA improve downstream task performance and robustness. H3: Given an initial amount of static robot manipulation data, it is more valuable to collect addition human mobile manipulation over mobile robot teleoperation data. We evaluate these hypotheses through three long horizon mobile manipulation tasks.
@@
Push Chair. Grasp the backrest of a office chair and push it against a table. This task requires coupled arm-based motion to move and align the orientation of the chair. We establish a base EMMA model by co-training with 10 min of mobile robot data and 20 min of human demonstrations. We use this to: (1) Compare EMMA’s performance to Mobile ALOHA trained on 30 min (10+20) of pure robot data. (2) Compare our unified EMMA policy against an ablation using separate navigation/manipulation models.
@@
Figure 5: (a) For Handover Wine task, starting with a fixed amount of static manipulation data, we show that adding more human fullbody motion data for EMMA (blue) yields to greater performance gains compared to adding mobile robot teleoperation data collected under an equivalent amount of time for Mobile ALOHA (orange). The performance gap expands from 10% to 30% as data increases from 15 to 60 minutes. (b) EMMA generates to an unseen scene with 54% full task success rate.
Evaluation protocols. We conduct 50 trials per task/model variant, recording error bars with 95% Clopper-Pearson confidence interval with 50 trials. To compare the success rates between EMMA and our baseline, we use a two-proportion Z-test. We claim statistically significant outperformance for a model if p << 0.05. Each trial is limited to 2 minutes. Any collisions are recorded as a failure for the corresponding subtask phase. All experiments use identical environmental conditions and object positions. We conducted 1150 mobile manipulation rollout evaluations in total.
@@ V-A Main Results
Figure 8: Time-lapse of all five tasks (variants) showing the task flow. EMMA is capable of learning tasks from human-robot interaction to precise bimanual coordination.
EMMA achieves favorable performance compared to systems trained on teleoperated mobile robot data (H1). EMMA consistently outperforms teleoperation-based baselines across a range of mobile-manipulation tasks by leveraging egocentric human demonstrations (Fig. 4). This quantitative gap is reinforced by qualitative evidence, as the Mobile ALOHA model exhibits frequent critical failures (Figure 9): These include (a) navigation collisions, such as hitting the TV, and (b) flawed bimanual grasps, which result in incorrectly pushing the chair. In the Handover Wine task, with the same 50 demos of static robot data, replacing one hour of teleoperated mobile robot data with one hour of human mobile manipulation yields an 82% success rate—a 30% increase over teleoperated robot data, which had a 52% success rate. This improvement demonstrates the benefit of learning from egocentric human-human interaction. Similarly, for the Push Chair task that requires coupled arm-base movements, adding 20 min of human data to 10 min of mobile robot data allows EMMA to match the performance of Mobile ALOHA trained on 30 min of pure robot teleoperation, (Figure 7). Leveraging egocentric human mobile manipulation data has also shown improvement in bimanual tasks found in the Grocery Shopping task. Lastly, using egocentric human mobile manipulation data has shown comparable performance to teleoperated mobile robot data on long-horizon Table Service task, a task that requires long-distance point-to-point navigation, where robot-teleoperated methods benefit from zero perspective gap. Taken together, these results show that egocentric human mobile manipulation data matches or even improves the value of equivalent robot teleoperation data and delivers safer, more reliable, and more scalable policies across diverse mobile-manipulation challenges.
EMMA generalizes to unseen scenes. We evaluate scene generalization by testing EMMA in a novel environment seen only through human demonstrations. Specifically, we collect 30 minutes of human demonstrations in a new spatial layout where the wine recipient stands randomly within a larger 5m $\times$ 2m area. Despite never seeing robot data from this environment, EMMA achieves 54% success rate (Fig. 5 (b)), demonstrating two key capabilities: (1) navigation behaviors learned from human data naturally adapt to the expanded spatial configuration, and (2) manipulation skills remain robust to visual domain shifts from the new environment. In contrast, Mobile ALOHA—trained exclusively on teleoperated data from the original environment is unable to complete the initial grasp due to the environmental changes (Fig. 5 (b)). This experiment validates that human data provides superior generalization through its inherent diversity compared to lab-constrained teleoperation.
@@
Ablation Studies (H2). Our ablation experiments on the Handover Wine task (Fig. 6 (a)) reveal the critical importance of both key components. Without kinematic retargeting, the success rate drops by 30%, as the robot frequently loses track of the human recipient when executing kinematically infeasible trajectories. The phase identification mechanism proves equally crucial—its removal causes complete task failure due to unintended base movements during grasping and unnecessary arm motions during navigation, resulting in collisions and dropped objects. These results validate that both human-robot embodiment alignment and learned phase modulation are essential for successful human-to-robot skill transfer in mobile manipulation. On the Push Chair task, a unified EMMA policy (10 min robot + 20 min human) outperforms separate navigation and manipulation models on precise bimanual chair grasp and continuous engagement. (Figure 7), indicating that single-policy co-training better exploits human data for arm-base coupling. These results validate that our design choices are essential for successful human-to-robot skill transfer. Our observed representative failure modes further reinforce this (Figure 9): (c) without phase switching, Mobile ALOHA makes unintended base motions that hinder precise manipulation. (d) Without kinematic retargeting, executing raw human waypoints yields infeasible trajectories;
@@
Human data scales mobile manipulation performance more efficiently compared to teleoperated robot data (H3). Keeping one hour of static robot manipulation data fixed, EMMA’s success rate climbs steadily from 0.36 to 0.82 as we increase human mobile manipulation data from 15 minutes to one hour. Mobile ALOHA also improves—from 0.26 to 0.52—when its robot teleoperation data grows from 15 minutes to one hour, but remains consistently below EMMA. This demonstrates that additional human mobile manipulation demonstrations yield substantially greater returns than equivalent increases in mobile robot teleoperation data.
VI Limitation

### v3

@@ Abstract
Scaling mobile manipulation imitation learning is bottlenecked by expensive mobile robot teleoperation. We present Egocentric Mobile MAnipulation (EMMA), an end-to-end framework training mobile manipulation policies from human mobile manipulation data with static robot data, sidestepping mobile teleoperation. To accomplish this, we co-train human full-body motion data with static robot data. In our experiments across four real-world tasks, EMMA demonstrates comparable performance to baselines trained on teleoperated mobile robot data (Mobile ALOHA), achieving higher or equivalent task performance in full task success. We find that EMMA is able to generalize to new spatial configurations and scenes, and we observe positive performance scaling as we increase the hours of human data, opening new avenues for scalable robotic learning in real-world environments. Details of this project can be found at: https://ego-moma.github.io
@@ I Introduction
Concurrently, recent work in cross-embodiment learning has explored human data as a scalable source of data for training end-to-end robot policies. Critically, human video data is cheap to collect, does not require a physical robot, and can be collected ergonomically via XR wearables [egomimic, qiu2025humanoidpolicyhuman]. Human video data has been leveraged to train better visual representations [nair2022r3muniversalvisualrepresentation], extract high-level object affordance [wang2023mimicplay, xu2024flowcrossdomainmanipulationinterface, wen2024anypointtrajectorymodelingpolicy, bharadhwaj2024track2actpredictingpointtracks], or even as direct action supervision via co-training [egomimic, qiu2025humanoidpolicyhuman]. However, these works have predominantly focused on table-top manipulation. To this end, we introduce Egocentric Mobile Manipulation (EMMA), an end-to-end system to train mobile manipulation policies solely from robot static-manipulation and human mobile-manipulation data. While the robot data is captured via teleoperation, the human data is captured simply by wearing Project Aria glasses [engel2023projectarianewtool], thus avoiding costly mobile manipulation teleoperation.
EMMA represents a step toward a new data paradigm in robot learning, where we imbue a robot with new skills by combining a set of robot teleoperation data with a more diverse pool of human demonstrations. In this work, our robot teleoperation data contains no mobile manipulation demonstrations, and we show that this skill can be effectively transferred from human data. EMMA presents a full-stack framework consisting of 1) an action retargeting pipeline which translates human mobile manipulation actions to a bimanual robot with a differential-drive base, 2) a unified architecture designed for co-training on heterogeneous human and robot data, and 3) an auxiliary phase identification mechanism that modulates between navigation and manipulation modes during inference. Therefore, it can switch control modes between navigation and manipulation phases based on predictions, preventing unintended base drift during manipulation and out-of-distribution arm movements during navigation.
We evaluate EMMA across four real-world mobile manipulation tasks: table service, handover wine, grocery shopping and push chair. Our experiments reveal three key findings. First, EMMA achieves superior performance compared to baselines trained on teleoperated mobile robot data, showing that human demonstrations can replace costly mobile teleoperation. Second, we observe favorable scaling properties—each additional portion of human data yields greater performance than an equivalent portion of teleoperated robot data. Third, EMMA exhibits robust generalization, transferring navigation and manipulation skills to novel environments seen only in human demonstrations. These results open new avenues for scalable robotic learning in real-world environments.
@@ II Related Work
Learning for Mobile Manipulation.
Building on successes in static manipulation, recent works have explored learning-based approaches for mobile manipulation that include skill primitives [sun2022fully, wu2023m, wu2023tidybot, kuang2025skillblender], reinforcement learning with decomposed action spaces  [gu2022multi, yokoyama2023asc, xia2021relmogen, xiong2024adaptive], and whole-body control objectives [fu2023deep, yang2024harmonic, hu2023causal, liu2024opt2skill, ben2025homie, wang2024hypermotion]. Unlike these approaches, end-to-end imitation learning enables mapping raw pixel information to whole-body actions, showing promising results through large-scale training [kim24openvla, ahn2022can, brohan2022rt, shafiullah2023bringing, pi0.5, moo2023arxiv]. A critical challenge is to collect high-quality mobile manipulation demonstrations. Pioneering works [jiang2025behavior, fu2024humanplus] have developed new teleoperation systems to facilitate data collection, including tethered leader-follower system [fu2024humanplus], VR headsets [chen2024arcapcollectinghighqualityhuman, lu2025mobiletelevisionpredictivemotionpriors], motion-capture suits [cisneros2022team, dafarra2024icub3, wholebodygeomretarget], smartphone-based control [pmlr-v164-wong22a], kinesthetic teaching [MOMO-FORCE], and full-body teleoperation for humanoid robots [fu2024humanplus]. However, despite these advances, collecting high-quality mobile manipulation data for diverse scenarios at scale remains a challenge. We propose to leverage egocentric human data (data collected with first-person perspective) captured by wearable devices as an alternative for scaling up imitation learning for mobile manipulation.
Robot Learning from Human Data. Recent work has focused on two complementary themes—leveraging human videos to bootstrap robot learning and finetuning with reinforcement learning for robust policies. In manipulation, co-training on paired egocentric human and robot demonstrations has been shown to boost skill performance [egomimic], while zero- or few-shot transfer from human videos can be enabled by image inpainting and motion-track priors [lepert2025phantomtrainingrobotsrobots, ren2025motion]. Building on these, hierarchical planners extract latent action sequences from humans and distill them via retargeters into whole-body controllers [wang2023mimicplay, fu2024humanplus, ImitationNet]. For navigation, policies become more reliable when imitation is fused with RL, either by converting keypoint matches into rewards or through behavior cloning bootstrapping [karnan2022voila, ramrakhya2023pirlnav]. Furthermore, inverse-dynamics models can pseudo-label passive egocentric video to distill intent-conditioned affordance subroutines [navsubroutine_egocentricvideos]. We aim to train mobile manipulation policies from human data in a unified learning framework.
@@ III Hardware and Data Preliminaries
Human and Robot Data Streams. We collect distinct data streams from both human demonstrations ($\mathcal{D}_{H}$) and robot execution ($\mathcal{D}_{R}$). A key shared data stream is the egocentric RGB image $I_{\text{ego}}$, generated by the Aria glasses worn by the human or mounted on the robot. The robot provides RGB streams from its two wrist cameras, $I_{\text{wrist}}$. Proprioceptive data streams capture the state of each embodiment: For the human, we leverage the Aria Machine Perception Service (MPS) [engel2023aria] to estimate bimanual 3D hand poses $H_{p}\in SE(3)\times SE(3)$ and the 3D head pose $H_{d}\in SE(3)$ (in a SLAM-based world frame). For the robot, proprioception includes the state of its bimanual arms via joint positions $R_{q}\in\mathbb{R}^{2\times 7}$ (including gripper state) and the corresponding end-effector poses $R_{p}\in SE(3)\times SE(3)$. These raw streams serve as the inputs to our system; their processing and transformation into policy inputs and action labels is a core algorithmic contribution detailed in Sec. IV-A.
@@ IV EMMA: System and Algorithm
EMMA is a scalable full-stack system, which enables (1) direct transfer of navigation skills from egocentric human data to a differential-drive mobile manipulator and (2) scaling up full mobile manipulation policy performance by co-training on both human mobile manipulation data and robot static manipulation data (Fig. 2).
@@ IV-B Human and Robot Data Co-training
By retargeting human demonstrations into executable robot trajectories (Sec. IV-A), we bridge the embodiment gap, allowing the co-training loop to treat heterogeneous human ($D_{H}$) and robot ($D_{R}$) data as a single, semantically aligned stream for the shared Transformer trunk.
Inspired by recent works in cross-embodiment policy learning [wang2024hpt, liu2024rdt, ghosh2024octo], we design an architecture based on a decoder-only Transformer with modality-specific input stems, a shared trunk, and multiple action and auxiliary output heads (Fig. 2).
Our goal is to transfer semantics of navigation and manipulation phases from human data. In addition, we want both the manipulation action prediction and navigation action prediction to benefit from large-scale human data. This motivates an architecture that shares a majority of the learnable parameters between the human and robot embodiments to promote shared representation, inspired by EgoMimic [egomimic].
@@
Co-training from heterogeneous sources. The model is trained jointly on batches drawn from two sources: (1) the collected human mobile manipulation demonstrations $D_{H}$ (with navigation actions processed by the retargeting module described in Sec. IV-A) and (2) static robot manipulation demonstrations $D_{R}$ (e.g., collected via teleoperation). When processing a human data batch, the human proprioception, shared ego vision stem, human manipulation action head, navigation action head and the shared phase prediction head are active. For a robot batch, the robot proprioception, wrist image stem, shared ego vision stem, and robot manipulation action head are active. The shared trunk and ego vision stem are updated by all sources and modalities, forcing them to learn versatile representations. The navigation head primarily learns from the retargeted human data, transferring human navigation strategies to the robot. The complete architecture is illustrated in Fig. 2.
@@ IV-C Auxiliary Phase Identification and Control Modulation
Phase Detection. We identify phases based on motion dynamics. For each frame, we compute the ratio of hand velocity to head velocity $r$ as detailed in Algorithm 1. To assess the accuracy of this unsupervised method, we hand-annotate 10% of uniformly sampled continuous segments from four tasks that require phase identification. We report Mean over Frames (MoF), defined as the number of frames where the manual label matches the dataset label divided by the total number of sampled frames.
@@
We empirically selected $\tau_{ratio}=2.0$, $\tau_{head}=0.4m/s$, and $\tau_{duration}=30$ based on the Table Service task. Consistently high MoF scores across tasks $(all\geq 0.92)$ confirm the system’s robustness to these specific threshold choices. To spatially localize $K$ manipulation zones, we fit a Gaussian Mixture Model (GMM) with $K=2$ components to head positions during these high-ratio periods. Each frame is classified as manipulation (phase 0) if its probability density under the GMM exceeds $\tau_{pdf}$, otherwise it is labeled as navigation (phase 1). GMM provides direct probability density estimates needed for classification without requiring labeled training data, while being computationally efficient and interpretable for spatial clustering.
@@
This phase-aware modulation removes the base action noise introduced from head movements during manipulation and ensures a smooth switch between the two phases.
@@ V Experiments
We aim to validate three key hypotheses. H1: EMMA can achieve performance comparable to systems trained on teleoperated mobile manipulation data. H2: Key design decisions of EMMA improve downstream task performance and robustness. H3: Given an initial amount of static robot manipulation data, it is more valuable to collect additional human mobile manipulation over mobile robot teleoperation data. We evaluate these hypotheses through four long horizon mobile manipulation tasks (Fig. 4).
Figure 4: Time-lapse of all five tasks (variants) showing the task flow. EMMA is capable of learning tasks from human-robot interaction to precise bimanual coordination. Numbers indicate chronological sequence.
@@
Push Chair. Grasp the backrest of an office chair and push it against a table. This task requires coupled arm-based motion to move and align the orientation of the chair. Unlike sequential tasks, EMMA here operates without phase-switching to demonstrate capacity for unified full-body coordination. We co-train a base EMMA model using 10 min of mobile robot data and 20 min of human demonstrations to: (1) compare against Mobile ALOHA trained on 30 min of robot-only data and (2) compare our unified policy against an ablation of separate navigation/manipulation models.
@@
Evaluation protocols. We conduct 50 trials per task/model variant, recording error bars with 95% Clopper-Pearson confidence interval. To compare the success rates between EMMA and our baseline, we use a two-proportion Z-test. We claim statistically significant outperformance for a model if p << 0.05. Each trial is limited to 2 minutes. Any collisions are recorded as a failure for the corresponding subtask phase. All experiments use identical environmental conditions and object positions. We conducted 1150 mobile manipulation rollout evaluations in total.
@@ V-A Main Results
EMMA achieves favorable performance compared to systems trained on teleoperated mobile robot data (H1). EMMA consistently outperforms teleoperation-based baselines across a range of mobile-manipulation tasks by leveraging egocentric human demonstrations (Fig. 5). This quantitative gap is reinforced by qualitative evidence, as the Mobile ALOHA model exhibits frequent critical failures (Fig. 6): These include (a) navigation collisions, such as hitting the TV, and (b) flawed bimanual grasps, which result in incorrectly pushing the chair. In the Handover Wine task, with the same 50 demos of static robot data, replacing one hour of teleoperated mobile robot data with one hour of human mobile manipulation yields an 82% success rate—a 30% increase over teleoperated robot data, which had a 52% success rate. This improvement demonstrates the benefit of learning from egocentric human-human interaction. Similarly, for the Push Chair task that requires coupled arm-base movements, adding 20 min of human data to 10 min of mobile robot data allows EMMA to match the performance of Mobile ALOHA trained on 30 min of pure robot teleoperation, (Fig. 7). Leveraging egocentric human mobile manipulation data has also shown improvement in bimanual tasks found in the Grocery Shopping task. Lastly, using egocentric human mobile manipulation data has shown comparable performance to teleoperated mobile robot data on long-horizon Table Service task, a task that requires long-distance point-to-point navigation, where robot-teleoperated methods benefit from zero perspective gap. Taken together, these results show that egocentric human mobile manipulation data matches or even improves the value of equivalent robot teleoperation data and delivers safer, more reliable, and more scalable policies across diverse mobile-manipulation challenges.
Figure 8: (a) EMMA generalizes to an unseen scene with 54% full task success rate. (b) For Handover Wine task, starting with a fixed amount of static manipulation data, we show that adding more human full-body motion data for EMMA (blue) yields greater performance gains compared to adding mobile robot teleoperation data collected in an equivalent amount of time for Mobile ALOHA (orange). The performance gap expands from 10% to 30% as data increases from 15 to 60 minutes.
EMMA generalizes to unseen scenes. We evaluate scene generalization by testing EMMA in a novel environment seen only through human demonstrations. Specifically, we collect 30 minutes of human demonstrations in a new spatial layout where the wine recipient stands randomly within a larger 5m $\times$ 2m area. Despite never seeing robot data from this environment, EMMA achieves 54% success rate (Fig. 8 (a)), demonstrating two key capabilities: (1) navigation behaviors learned from human data naturally adapt to the expanded spatial configuration, and (2) manipulation skills remain robust to visual domain shifts from the new environment. In contrast, Mobile ALOHA trained exclusively on teleoperated data from the original environment is unable to complete the initial grasp due to the environmental changes (Fig. 8 (a)). This experiment validates that human data provides superior generalization through its inherent diversity compared to lab-constrained teleoperation.
@@
Ablation Studies (H2). Our ablation experiments on the Handover Wine task (Fig. 9) reveal the critical importance of both key components. Without kinematic retargeting, the success rate drops by 30%, as the robot frequently loses track of the human recipient when executing kinematically infeasible trajectories. The phase identification mechanism proves equally crucial—its removal causes complete task failure due to unintended base movements during grasping and unnecessary arm motions during navigation, resulting in collisions and dropped objects. These results validate that both human-robot embodiment alignment and learned phase modulation are essential for successful human-to-robot skill transfer in mobile manipulation. On the Push Chair task, a unified EMMA policy (10 min robot + 20 min human) outperforms separate navigation and manipulation models on precise bimanual chair grasp and continuous engagement. (Fig. 7), indicating that single-policy co-training better exploits human data for arm-base coupling. These results validate that our design choices are essential for successful human-to-robot skill transfer. Our observed representative failure modes further reinforce this (Fig. 6): (c) without phase switching, Mobile ALOHA makes unintended base motions that hinder precise manipulation. (d) Without kinematic retargeting, executing raw human waypoints yields infeasible trajectories.
@@
Human data scales mobile manipulation performance more efficiently compared to teleoperated robot data (H3). Keeping one hour of static robot manipulation data fixed, EMMA’s success rate climbs steadily from 0.36 to 0.82 as we increase human mobile manipulation data from 15 minutes to one hour (Fig. 8 (b)). Mobile ALOHA also improves from 0.26 to 0.52 when its robot teleoperation data grows from 15 minutes to one hour, but remains consistently below EMMA. This demonstrates that additional human mobile manipulation demonstrations yield substantially greater returns than equivalent increases in mobile robot teleoperation data.
VI Limitations